OpenStudy (samsan9):

Describe the end behavior of the graph for each function. f(x)=x^3+2x^2-5x+1

OpenStudy (anonymous):

End behavior can be described by the limits\[\lim_{x\to\pm\infty}f(x)\]

OpenStudy (samsan9):

what about g(x)=-2x^3-8x^2+18x+72=0

OpenStudy (samsan9):

@SithsAndGiggles

OpenStudy (anonymous):

The procedure is the same. For example, directly substituting, you have \[\lim_{x\to-\infty}g(x)=-2(-\infty)^3-8(-\infty)^2+18(\infty)+72=+\infty_1-\infty_2+\infty_3\] Now, \(\infty-\infty\) isn't properly defined. This is why I used different subscripts to denote different \(\infty\)s. In this case, since \(x^3\) increases faster than \(x^2\), the \(\infty\) from the first term \((-2x^3)\) dominates the \(\infty\) from the second term \((-8x^2)\), which means that \(\infty_1>\infty_2\), so \(\infty_1-\infty_2=\infty\). So, the limit as \(x\to-\infty \) of \(g(x)\) is \(+\infty\). Does this make sense?

OpenStudy (samsan9):

well ill take a picture and read through it to make sure :) thank you

OpenStudy (anonymous):

np